For many years, Thomson Company manufactured a single product called LEC 40. Then three years ago, the company automated a portion of its plant and at the same time introduced a second product called LEC 90 that has become increasingly popular. The LEC 90 is a more complex product, requiring 0.80 hours of direct labor time per unit to manufacture and extensive machining in the automated portion of the plant. The LEC 40 requires only 0.60 hours of direct labor time per unit and only a small amount of machining. Manufacturing overhead costs are currently assigned to products on the basis of direct labor-hours.

Despite the growing popularity of the company’s new LEC 90, profits have been declining steadily. Management is beginning to believe that there may be a problem with the company’s costing system. Direct material and direct labor costs per unit are as follows:

LEC 40 LEC 90
Direct materials $ 30.00 $ 50.00
Direct labor (0.60 hours and 0.80 hours @ $10.00 per hour) $ 6.00 $ 8.00

Management estimates that the company will incur $960,000 in manufacturing overhead costs during the current year and 50,000 units of the LEC 40 and 25,000 units of the LEC 90 will be produced and sold.

Required:

1-a. Compute the predetermined overhead rate assuming that the company continues to apply manufacturing overhead cost on the basis of direct labor-hours.

1-b. Using this rate and other data from the problem, determine the unit product cost of each product.
Answer #1

1-a. Pre-determined overhead rate = Total estimated manufacturing overhead / Total estimated direct labor hours

Total estimated direct labor hours = Labor hours to produce 50,000 units LEC 40 + Labor hours to produce 25,000 units of LEC 90 = ( 0.60 hours * 50,000 ) + ( 0.80 hours * 25,000 ) = 50,000 hours

Pre-determined overhead rate = $960,000 / 50,000 hours = $19.20 per hour.

1-b. unit product cost of each product

Unit product cost of LEC 40 = Direct materials cost per unit + Direct labor cost per unit + manufacturing cost per unit

Manufacturing cost per unit of LEC 40 = Pre-determined overhead rate * Labor hours per unit = $19.20 * 0.6 hours = $11.52

Unit product cost of LEC 40 = $30 + $6 + $11.52 = $47.52 per unit.

Unit product cost of LEC (0 = Direct materials cost per unit + Direct labor cost per unit + manufacturing cost per unit

Manufacturing cost per unit of LEC 90 = Pre-determined overhead rate * Labor hours per unit = $19.20 * 0.8 hours = $15.36

Unit product cost of LEC 90 = $50 + $8 + $15.36 = $73.36 per unit.